OpenStudy (calculator):

✰✰✰✰★Help me derive this equation★✰✰✰✰ At point O on its path, a projectile has speed V and is travelling at an angle α above the horizontal. Derive the equation of the trajectory of the projectile in the form below

OpenStudy (calculator):

\[y=xtan(\alpha)-\frac{gx^2}{2V^2}(1+\tan^2(\alpha))\] where x- and y- axes pass through O and are directed horizontally and vertically upwards respectively.

OpenStudy (calculator):

Although I'm good at projectile problems, but this messes me up. How do you start?

OpenStudy (anonymous):

The left side:\[x=v_{x}t\] \[v _{x}=vcos \alpha\] \[xtan \alpha=v (\cos \alpha )t(\sin \alpha)/\cos \alpha\] \[=vtsin \alpha=v _{y}\]

OpenStudy (anonymous):

*v(y)t

OpenStudy (calculator):

I don't even understand what you're doing

OpenStudy (anonymous):

 x=v(x)t as distance= speed*time The third line is replacing in x with v(x)t=vcos(a)t tanx=sinx/cosx The 3rd line simplifies to the 4th line, and [vsin(a)]t=[v(y)]t=y The right side thing is with gravity added in

OpenStudy (anonymous):

The right side: acceleration(y)= -g v(y)=-gt y=-0.5g t^2 This is from calculus. Do you understand this?

OpenStudy (anonymous):

Anyway, adapting the last formula to suit your question; \[t^2=\frac{x^2+x^2 \tan^2( \alpha)}{v^2}\] \[t^2 v^2=x^2(1+ \tan^2( \alpha))\] \[tv=\sqrt{x^2(1+\tan^2 (\alpha))}= \sqrt{x^2}\sqrt{1+ \tan^2(\alpha)}\] \[tv=x \sqrt{1+\tan^2 \alpha}\]

OpenStudy (anonymous):

tana=v(y)/v(x) (as tanx= opposite/adjacent side)

OpenStudy (anonymous):

\[\frac{tv}{x}= \sqrt{1+ \frac{v _{y}^2}{v _{x}^2}}\] multiplying the right side by v(x)/(v(x)...

OpenStudy (calculator):

\[u_y=usinx \\ \\ u_x=ucosx\] \[u_x=v_x \\ \\ u_x=\frac{x}{t}\] \[t=\frac{x}{u_x}\] From equation \[y=u_yt-\frac{1}{2}g t^2\] \[y=(usinx)t-\frac{1}{2}g t^2\] \[y=usin(x)(\frac{x}{u_x})-\frac{1}{2}g(\frac{x^2}{u_x^2}) \\ \\ y=\cancel{u}\sin(x)(\frac{x}{\cancel{u}\cos(x)})-\frac{1}{2}g\frac{x^2}{u^2\cos^2x} \\ \\ y=xtan(x)-\frac{1}{2}g\frac{x^2}{(u^2\cos^2x)}\] I got until here

OpenStudy (anonymous):

\[tv/x=v/v _{x}\] \[t=x/v _{x}\] It seems that I've been working backwards- but trace back my steps and you'll get to that equation.

OpenStudy (calculator):

I think i got it \[y=xtan(\alpha)-\frac{gx^2}{2u^2\cos^2\alpha}\] \[y=xtan(\alpha)-\frac{gx^2}{2u^2}(sec^2(\alpha))\] \[y=xtan(\alpha)-\frac{gx^2}{2u^2}(1+tan^2(\alpha))\] YES

OpenStudy (anonymous):

I'm glad you got it in spite of my cryptic answering.

OpenStudy (calculator):

Thanks to walter lewin too :D http://www.youtube.com/watch?v=0OV1E9M9Kkc

OpenStudy (anonymous):

Odd how omni-pervasive he is: I was considering offering that link by means of an answer..